A spherical drop of mercury of radius R has a capacitance given by Cm = 4piepsilonoR. If 8 such drops combine to form a single larger drop, what is its capacitance? Give your answer in terms of Cm
Answer #1

Given : n = 8 ; Cm = 40R

Solution :

As given ,8 drops combine to form a single large drop.

Volume of big drop = 8 times the volume of a small drop

Volume of sphere = (4/3)R3

Since volume is proportional to the radius cubed . Hence the radius of combined drop is given by : R = 81/3 = 2 times that of smaller drop.

As capacitance is directly proportional to the radius. The capacitance will also be increased by the same factor.

i.e. C = 2Cm

Answer : C = 2 Cm